# Multi-granularity Interest Retrieval and Refinement Network for Long-Term User Behavior Modeling in CTR Prediction

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709438isbn: 979-8-4007-1245-6/25/08ccs: Information systemsccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Click-Through Rate (CTR) prediction is a crucial task in modern online personalization platforms (Wang et al., 2021a, 2019, b; Zhang et al., 2024; Wang et al., 2024b), such as recommendation systems (Tong et al., 2024; Wu et al., 2024; Yin et al., 2024c) and online advertising (Estrada-Jiménez et al., 2017; Brajnik and Gabrielli, 2010). Its objective is to estimate the probability of a user clicking on a specific item within a given contextual situation (Chen et al., 2016; Guo et al., 2017). Recent advancements in CTR prediction have shown that user behavior modeling which incorporates the user’s historical behavior sequence to learn implicit interests results in significant performance gain (Zhou et al., 2018a, b, 2019). As online platforms evolve, the length of user behavior sequences has experienced a surge.For instance, on Taobao, one of China’s largest e-commerce platforms, over 23% of users have recorded more than 1,000 clicks in the past five months alone (Ren et al., 2019). Effectively modeling this massive and intricate user behavior data has emerged as a focal point for both academia and industry.
Due to the strict online inference time constraints, previous methods (Zhou et al., 2018b, 2019; Kang and McAuley, 2018) considering their high time complexity, can only utilize the most recent behaviors with the sequence length seldom exceeding 100. This limitation leads to suboptimal results. To address this, several models have been proposed to leverage long behavior sequences for modeling user interests. Most of these models follow two cascading stages (Pi et al., 2020). In the first stage, the General Search Unit (GSU) identifies a small number of behaviors most relevant to the target item from thousands of long-term behaviors using fast and coarse search methods. In the second stage, the Exact Search Unit (ESU) captures the relationships between behaviors in these subsequences and the target item to precisely learn user interests. The most critical difference among these models such as SIM (Pi et al., 2020), UBR4CTR (Qin et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), UBCS (Zhang et al., 2022) and TWIN (Chang et al., 2023), lies in the different search strategies adopted by GSU.
@@
Figure 1. The left side of the picture illustrates a user’s behavior sequence, with black sneakers as the target item. The middle column illustrates the microcosm of users’ interests at different granularities, which represents the multi-granularity queries we aim to construct. The right column displays the subsequences related to the multi-granularity interests.
@@ 2.1. Click-Through-Rate Prediction
Early CTR models primarily concentrate on exploiting feature interactions. Pioneering studies are relatively shallow such as factorization machines(FM) (Rendle, 2010) and its variants (Pan et al., 2016; Juan et al., 2016; Pan et al., 2018). With the great success of deep learning, various studies such as Wide&Deep (Cheng et al., 2016), PNN (Qu et al., 2016), DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017, 2021c), successfully combine deep and shallow models. Moreover, several models (Xiao et al., 2017; He and Chua, 2017; Lian et al., 2018; Yu et al., 2023, 2024) have explored more complex neural networks, such as xDeepFM (Lian et al., 2018) further introduces compressed interaction networks to explicitly model high-order feature interactions.
With the increasing demand for personalization, researchers are focusing on user behavior modeling (Zhou et al., 2018b; Han et al., 2023; Yin et al., 2023; Shen et al., 2024; Zhou et al., 2019, 2024), which derives the latent representation of users’ interest from their behavior sequences. YoutubeDNN (Covington et al., 2016) performs mean pooling on the entire user sequence. DIN (Zhou et al., 2018b) proposes an attention mechanism to identify target-relevant interests. DIEN (Zhou et al., 2019) introduces a novel GRU (Cho et al., 2014) with an attention update gate (AUGRU) to capture the evolving process of user interests. DSIN (Feng et al., 2019) observes that user behaviors are highly homogeneous in each session, and heterogeneous cross sessions. DMIN (Xiao et al., 2020) and MIND (Li et al., 2019) employ self-attention mechanisms and capsule networks, respectively, to represent user interests as multiple vectors. CAN (Zhou et al., 2020) captures the feature co-action by directing user behavior features into a compact multi-layer perceptron generated by target features.
@@ 2.2. Long-Term User Behavior Modeling
As user behavior modeling (Yin et al., 2024a; Han et al., 2024; Wang et al., 2024a; Xie et al., 2024a; Yin et al., 2024b; Xie et al., 2024b; Zhou et al., 2018b) has demonstrated remarkable performance in industrial applications, researchers start to focus on modeling increasingly longer behaviors. Due to strict limitations on inference time and model parameter count, DIN-like models (Zhou et al., 2018b, 2019; Feng et al., 2019) are less suitable for long sequence modeling.
@@ 4.2. Overall Performance
(1) Our proposed MIRRN consistently outperforms all baselines across three public datasets, illustrating its effectiveness. Specifically, MIRRN achieves AUC improvements of 0.67%, 1.35%, and 0.51% improvements over the strongest baseline on the Taobao, Alipay and Tmall datasets, respectively.
(2) The top two strongest baselines on the three datasets predominantly come from models on long-term behavior modeling. Among the three types of baselines, long-term behavior modeling models generally deliver the best performance, followed by short-term behavior modeling models. The experiments demonstrate that modeling user interest significantly enhances CTR model performance. Furthermore, leveraging more extensive user historical behaviors can lead to substantial performance gains.
(3) Among the short-term behavior modeling models, DMIN achieved the best performance on two datasets. DMIN emphasizes that users typically have multiple interests. It uses multi-head self-attention to implicitly extract multiple user interests. The strong performance of DMIN highlights the effectiveness of learning users’ multiple interests, which is similar to the motivation behind our model. Our proposed MIRRN further enhances performance by efficiently utilizing users’ long-term behaviors and explicitly capturing diverse interests, thereby surpassing DMIN.
(4) Among the long-term behavior modeling baselines, SIM Soft and TWIN achieve strong results. SIM Soft uses the dot product of pre-trained embeddings as the relevance metric, while TWIN maintains consistency in relevance metrics during the two-stage interest modeling. However, both methods focus solely on the user’s target-relevant behaviors during the first stage and lack the mining of relational information within subsequences during the second stage. In contrast, MIRRN considers the user’s multi-granularity interests during the retrieval stage and effectively mines sequential and interactive information within subsequences, resulting in superior performance. These points will be further verified in the subsequent ablation study.
@@ 4.3. Ablation Study
(1)We assess the effectiveness of capturing multi-granularity interests by evaluating the three search units in MIRM, which are denoted as TASU, LASU and GASU. The experimental results, presented in Table 4, are obtained with a constant total number of retrieved behaviors. Notably, the model using only TASU performs relatively well, which may explain why previous models focusing solely on target-aware interest still achieved notable results. In contrast, models using only LASU or GASU exhibit poor performance. Moreover, any model lacking TASU experiences a significant performance decline, underscoring the critical importance of TASU. Nevertheless, adding any of the search units consistently enhances performance, demonstrating the value of capturing users’ diverse interests. Our proposed MIRRN, which employs all three search units, effectively captures interests at different granularities, leading to the best performance.
(2)We validate the effectiveness of BSRM and MIAM by separately removing each component. The results, as shown in Table 5, reveal that removing either BSRM or MIAM results in a performance decrease. It indicates that both mining the relational information of behaviors and adaptively learning the impact of multi-granularity interests on the target item are crucial.
(3)We further verify the effectiveness of two key components, TAPE and MHFT, in BSRM, which are responsible for mining sequential and interactive information respectively. The results in Table 5 demonstrate that both sequential information in TAPE and interactive information in MHFT positively influence performance.
(4) We conduct comparative experiments on three methods for mining interactive information in sequences: FT, MHFT and MHSA. The results in Table 5 indicate that our proposed MHFT achieves performance comparable to MHSA and consistently outperforms FT. MHFT even achieves the best results on two datasets. Notably, MHFT exhibits lower time complexity and requires fewer parameters as explained in Section 3.6.2. Therefore, MHFT demonstrates superiority among these three methods.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709438isbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systems
@@ 1. Introduction
Click-Through Rate (CTR) prediction is a crucial task in modern online personalization platforms, such as recommendation systems and online advertising (Estrada-Jiménez et al., 2017; Brajnik and Gabrielli, 2010). Its objective is to estimate the probability of a user clicking on a specific item within a given contextual situation (Chen et al., 2016; Guo et al., 2017). Recent advancements in CTR prediction have shown that user behavior modeling which incorporates the user’s historical behavior sequence to learn implicit interests results in significant performance gain (Zhou et al., 2018a, b, 2019). As online platforms evolve, the length of user behavior sequences has experienced a surge.For instance, on Taobao, one of China’s largest e-commerce platforms, over 23% of users have recorded more than 1,000 clicks in the past five months alone (Ren et al., 2019). Effectively modeling this massive and intricate user behavior data has emerged as a focal point for both academia and industry.
Due to the strict online inference time constraints, previous methods (Zhou et al., 2018b, 2019; Kang and McAuley, 2018; Sun et al., 2019) considering their high time complexity, can only utilize the most recent behaviors with the sequence length seldom exceeding 100. This limitation leads to suboptimal results. To address this, several models have been proposed to leverage long behavior sequences for modeling user interests. Most of these models follow two cascading stages (Pi et al., 2020). In the first stage, the General Search Unit (GSU) identifies a small number of behaviors most relevant to the target item from thousands of long-term behaviors using fast and coarse search methods. In the second stage, the Exact Search Unit (ESU) captures the relationships between behaviors in these subsequences and the target item to precisely learn user interests. The most critical difference among these models such as SIM (Pi et al., 2020), UBR4CTR (Qin et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), UBCS (Zhang et al., 2022) and TWIN (Chang et al., 2023), lies in the different search strategies adopted by GSU.
@@
Figure 1. The left side of the picture depicts a user’s behavior sequence, with black sneakers as the target item. The middle column represents the microcosms of the user’s multi-granularity interests, corresponding to the diverse queries we aim to construct. The right column displays the subsequences related to the multi-granularity interests.
@@ 2.1. Click-Through-Rate Prediction
Early CTR models primarily concentrate on exploiting feature interactions. Pioneering studies are relatively shallow such as factorization machines(FM) (Rendle, 2010) and its variants (Pan et al., 2016; Juan et al., 2016; Pan et al., 2018). With the great success of deep learning, various studies such as Wide&Deep (Cheng et al., 2016), PNN (Qu et al., 2016), DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017, 2021), successfully combine deep and shallow models. Moreover, several models (Xiao et al., 2017; He and Chua, 2017; Lian et al., 2018; Yu et al., 2023, 2024) have explored more complex neural networks, such as xDeepFM (Lian et al., 2018) further introduces compressed interaction networks to explicitly model high-order feature interactions.
With the increasing demand for personalization, researchers are focusing on user behavior modeling (Zhou et al., 2018b; Yin et al., 2023; Shen et al., 2024; Zhou et al., 2019, 2024), which derives the latent representation of users’ interest from their behavior sequences. YoutubeDNN (Covington et al., 2016) performs mean pooling on the entire user sequence. DIN (Zhou et al., 2018b) proposes an attention mechanism to identify target-relevant interests. DIEN (Zhou et al., 2019) introduces a novel GRU (Cho et al., 2014) with an attention update gate (AUGRU) to capture the evolving process of user interests. DSIN (Feng et al., 2019) observes that user behaviors are highly homogeneous in each session, and heterogeneous cross sessions. DMIN (Xiao et al., 2020) and MIND (Li et al., 2019) employ self-attention mechanisms and capsule networks, respectively, to represent user interests as multiple vectors. CAN (Zhou et al., 2020) captures the feature co-action by directing user behavior features into a compact multi-layer perceptron generated by target features.
@@ 2.2. Long-Term User Behavior Modeling
As user behavior modeling (Han et al., 2024; Yin et al., 2024; Xie et al., 2024; Zhou et al., 2018b) has demonstrated remarkable performance in industrial applications, researchers start to focus on modeling increasingly longer behaviors. Due to strict limitations on inference time and model parameter count, DIN-like models (Zhou et al., 2018b, 2019; Feng et al., 2019) are less suitable for long sequence modeling.
@@ 4.2. Overall Performance
(1). Our proposed MIRRN consistently outperforms all baselines across three public datasets, illustrating its effectiveness. Specifically, MIRRN achieves AUC improvements of 0.67%, 1.35%, and 0.51% improvements over the strongest baseline on the Taobao, Alipay and Tmall datasets, respectively.
(2). The top two strongest baselines on the three datasets predominantly come from models on long-term behavior modeling. Among the three types of baselines, long-term behavior modeling models generally deliver the best performance, followed by short-term behavior modeling models. The experiments demonstrate that modeling user interest significantly enhances CTR model performance. Furthermore, leveraging more extensive user historical behaviors can lead to substantial performance gains.
(3). Among the short-term behavior modeling models, DMIN achieved the best performance on two datasets. DMIN emphasizes that users typically have multiple interests. It uses multi-head self-attention to implicitly extract multiple user interests. The strong performance of DMIN highlights the effectiveness of learning users’ multiple interests, which is similar to the motivation behind our model. Our proposed MIRRN further enhances performance by efficiently utilizing users’ long-term behaviors and explicitly capturing diverse interests, thereby surpassing DMIN.
(4). Among the long-term behavior modeling baselines, SIM Soft and TWIN achieve strong results. SIM Soft uses the dot product of pre-trained embeddings as the relevance metric, while TWIN maintains consistency in relevance metrics during the two-stage interest modeling. However, both methods focus solely on the user’s target-relevant behaviors during the first stage and lack the mining of relational information within subsequences during the second stage. In contrast, MIRRN considers the user’s multi-granularity interests during the retrieval stage and effectively mines sequential and interactive information within subsequences, resulting in superior performance. These points will be further verified in the subsequent ablation study.
@@ 4.3. Ablation Study
(1). We assess the effectiveness of capturing multi-granularity interests by evaluating the three search units in MIRM, which are denoted as TASU, LASU and GASU. The experimental results, presented in Table 4, are obtained with a constant total number of retrieved behaviors. Notably, the model using only TASU performs relatively well, which may explain why previous models focusing solely on target-aware interest still achieved notable results. In contrast, models using only LASU or GASU exhibit poor performance. Moreover, any model lacking TASU experiences a significant performance decline, underscoring the critical importance of TASU. Nevertheless, adding any of the search units consistently enhances performance, demonstrating the value of capturing users’ diverse interests. Our proposed MIRRN, which employs all three search units, effectively captures interests at different granularities, leading to the best performance.
(2). We validate the effectiveness of BSRM and MIAM by separately removing each component. The results, as shown in Table 5, reveal that removing either BSRM or MIAM results in a performance decrease. It indicates that both mining the relational information of behaviors and adaptively learning the impact of multi-granularity interests on the target item are crucial.
(3). We further verify the effectiveness of two key components, TAPE and MHFT, in BSRM, which are responsible for mining sequential and interactive information respectively. The results in Table 5 demonstrate that both sequential information in TAPE and interactive information in MHFT positively influence performance.
(4). We conduct comparative experiments on three methods for mining interactive information in sequences: FT, MHFT and MHSA. The results in Table 5 indicate that our proposed MHFT achieves performance comparable to MHSA and consistently outperforms FT. MHFT even achieves the best results on two datasets. Notably, MHFT exhibits lower time complexity and requires fewer parameters as explained in Section 3.6.2. Therefore, MHFT demonstrates superiority among these three methods.